OpenStudy (anonymous):

Ratios and proportions. A ranger wants to estimate the number of antelope in a preserve. He catches 58 antelope, tags their ears, and returns them. Some time later, he catches 29 antelope and discovers that 7 of them are tagged. Estimate the number of antelope in the preserve.

OpenStudy (anonymous):

a/b = c/d

OpenStudy (anonymous):

can someone check this please... I took x/58 = 7/29 29x = 406 x = 14 58 + 14 = 72 roughly 72 antelope in the preserve????

OpenStudy (anonymous):

your variable x relates to the estimated total antelope in the preserve. Since he caught and tagged 58 out of x, I think you'll want to switch your left hand equation to 58/x instead of x/58. Then you have 58/x : 7/29

OpenStudy (anonymous):

7 into 1682 = 240.285....

OpenStudy (anonymous):

I put 58 on the bottom and 29 on the bottom to match those were caught

OpenStudy (anonymous):

I added 14 to 58 because its the difference from those caught and tagged

OpenStudy (anonymous):

Yes, but the fact that they are caught is not the association that you want for this ratio. He originally caught 58 of an undetermined number = ratio1. He then catches 29, of which 7 are tagged = ratio2. The reason you want the x paired with the 29 is that the 29 represents a sample set with which you want to estimate the total population.

OpenStudy (anonymous):

how do I answer the question of : estimate the number of antelope in the preserve... with that ratio? or solve it?

OpenStudy (anonymous):

Okay. So we have our ratio of 58 7 -- : -- x 29 which can be read as 58 is to 7 as x is to 29. You calculated the number above, by multiplying the top on one side by the bottom on the other: 58*29 = 1682 and then dividing by 7 = 240.285. So the estimate would be 240 antelope.

OpenStudy (anonymous):

makes sense... thank you